OpenStudy (fairy733189):

Which statement best describes the graph of -x3 - x2 + 4x + 4? t starts down on the left and goes up on the right and intersects the x-axis at x = -2, 2, and 4. It starts down on the left and goes up on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, 2, and 4.

OpenStudy (fairy733189):

@mathmale

OpenStudy (mathmale):

First: a bit of housekeeping. Your "-x3 - x2 + 4x + 4" needs to be written as f(x) = -x^3 - x^2 + 4x + 4, with the ^ symbol indicating exponentiation. First question (for you): What does the graph of y=x^3 look like? Could you draw it in the Draw utility, below?

OpenStudy (fairy733189):

I actually don't really know what it looks like, can you xplain that to me?

OpenStudy (mathmale):

It'd be very smart of you to look up basic functions such as y=x, y=x^2, y=x^3, and so on, along with their graphs, and to work at memorizing these basic graphs. You will see these again and again in the future.